# Maintaining Distributed Data Structures in Dynamic Peer-to-Peer Networks

### v1

@@ Appendix A Useful Mathematical Tools
for $\text{for }R\geq 6\mathbb{E}\left[X\right]$ (1)
@@
for $\text{for }k>\mu$ (2)
for $\text{for }\varepsilon\geq 0$ (3)
@@
for $\text{for }k<\mu$ (4)
for $\text{for }\varepsilon\in(0,1)$ (5)
@@ Appendix B Useful Skip list properties
The size of a run of nodes $X_{\ell}$ for some level $\ell$, is at most $\mathcal{O}(\log n)$ w.h.p., and its expected value is $1/(1-p)$.
@@
The size of a run of nodes $|X_{\ell}|$ is a geometric random variable with parameter $1-p$. Thus, Choosing $k=c\log_{1-p}n+1$ for $c\geq 1$, gives us a high confidence bound on the number of consecutive nodes in a run of node $X_{\ell}$ at some level $\ell$. Indeed, Moreover, its expected value is $\mathbb{E}\left[|X_{\ell}|\right]=1/(1-p)$, assuming $p=1/2$ (in other words, constant), we have that the expected length of a run of nodes at some level $\ell$ of the skip list is $\mathcal{O}(1)$.
@@
Let $R_{i}$ be number of horizontal edges at level $0\leq i\leq h$ crossed by a search operation that starts on the left topmost node of the skip list. Define the random variable $W_{h}=|R_{0}|+|R_{1}|+\dots+|R_{h}|$ to be the amount of horizontal moves performed by the search algorithm. Observe that each $R_{i}$ is a geometric random variable of parameter $1-p$ and that $h$ is a random variable itself. From Lemma B.1 we know that $\mathbb{E}\left[h\right]=\mathcal{O}(\log n)$ and that . Since $h$ is a random variable, $W_{h}$ is a random sum of random variables thus we can not make a straightforward use of the properties of sum of geometric random variables, rather we can write:
@@
Observe that the first sum is bounded above $16\log n$ because every probability is less than $1$ and the second one is dominated by $\sum_{i\geq 1}1/i^{2}$ which is a constant. Since the running time of a search operation is bounded by the number of horizontal moves performed at each level plus the number of vertical moves to reach the target node from the topmost level to the bottom most. We have that the overall running time of a search procedure starting at the left topmost node in the skip list is $T=h+W_{h}=\mathcal{O}(\log n)$ with probability at least $1-(1/n^{2})$, by applying the union bound we have that $T$ is $\mathcal{O}(\log n)$ with probability at least $1-(1/n)$ starting at any node in the skip list. To conclude, the expected running time is $\mathbb{E}\left[T\right]=\mathbb{E}\left[h\right]+\mathbb{E}\left[W_{h}\right]=\mathcal{O}(\log n)$.
@@ Appendix C Spartan’s Reshaping protocol
Figure 12: Example of the reshaping procedure. On the left, the Spartan structure as a one-dimensional wrapped butterfly of committees. The rows are represented with their binary representations. Every node in the figure is a random committee of $\Theta(\log n)$ nodes and every edge between two committees encodes a complete bipartite graph. On the right, $2$-dimensional butterfly is obtained by increasing the dimensionality of the left-wrapped butterfly by one. Green nodes are the ones that must be added to increase the dimensionality of the $1$-dimensional wrapped butterfly. Moreover, left-to-right execution increases the dimensionality by one, and right-to-left execution decreases the dimensionality by one.
@@
In this section, we provide a detailed description of the reshaping protocol mentioned in Section 2.1. Let $\alpha,\beta\in\mathbb{R}_{\geq 0}$ such that $\alpha>1$ and $\beta\in(0,1)$ be two constants. We want the Spartan network $\mathcal{S}$ to be enlarged if all the committees have size at most $\alpha c\log n$, to shrink if all the committees have size at most $\beta c\log n$, or stay the same otherwise ($c>0$ is the constant hidden by the asymptotic notation). In other words, we want $\mathcal{S}$ to be enlarged or reduced when the wrapped butterfly network size increases or decreases by one dimension respectively. Moreover, a general technique to increase the dimensionality of a $k$-dimensional butterfly network $\mathcal{S}$ by one is to: (1) create a copy of $\mathcal{S}$ in which each row ID $(i,j)$ is “shifted” by $2^{k}$ i.e., $(2^{k}+i,j)$; (2) “shift” the index of the levels of the two disjoint butterflies by one; (3) add a new row of $2^{k+1}$ nodes to form the level 00 of the $k+1$ dimensional butterfly; and, (5) connect the newly created level 00 to the two disjoint butterflies using the “standard” butterfly edge-construction rule [53]. Furthermore, to reduce the dimensionality of a $k$-dimensional butterfly, we need to remove the level 00, remove one of the two disjoint $k-1$-dimensional butterflies, and fix the row and levels IDs if needed. Figure 12 shows an example on how to perform such operations. We start with the description of the growing procedure. A new number of nodes $n^{\prime}=\alpha n$ is reached such that $2^{k}k\in\mathcal{O}(N)$ must be increased to $2^{k+1}(k+1)\in\mathcal{O}(N^{\prime})$ (the number of committees increases from $N=\mathcal{O}(n/\log n)$ to $N^{\prime}=\mathcal{O}(n^{\prime}/\log n^{\prime})$) and update the butterfly structure according to the new size of the network. Moreover, each step of Algorithm 3 requires at most $\mathcal{O}(\log n)$ rounds and no node will send or receive more than $\mathcal{O}(\log n)$ messages at any round. As a first step, the committee $C(0,0)$ is elected as a leader [^13] We can make this assumption because we do not deal with byzantine nodes., and each committee routes a message in which it expresses its opinion (grow, shrink, stay the same) to the leader committee (lines 1-2). This can be done very efficiently on hypercubic networks [53, 61], indeed we have that the leader committee will receive all the opinions in $\mathcal{O}(\log N)$ rounds [54] where $N\in\mathcal{O}(n/\log n)$ is the number of committees in the wrapped butterfly. If the network agrees on growing, the protocol proceeds with increasing the dimensionality of the wrapped butterfly by one. Moreover, assume that the $k$-dimensional butterfly must be transformed into a $k+1$-dimensional one. To this end, the protocol executes the growing approach described before. Every committee shifts its level ID by one and promotes two random nodes in each committee to be the committee leaders of the copy of the butterfly and of the new level 00. Next, all the newly elected committee leaders will start recruiting random nodes in the network until they reach a committee size of $\Theta(\log n^{\prime})$ where $n^{\prime}=\alpha n$. Finally, each new committee will create edges according to the wrapped butterfly construction algorithm [53, 13] and drop the old ones (if any). The next lemma shows that the growing phase requires $\mathcal{O}(\log n^{\prime})$ rounds w.h.p. where $n^{\prime}=\alpha n$.
@@
[…]
Furthermore, in the case in which the network agreed on shrinking i.e., reducing the butterfly dimensionality from $k$ to $k-1$ a pool of vacated nodes $\mathcal{V}$ is created and all the nodes in the level 00 and in the committees $C(i,j)$ such that $2^{k-1}\leq i<2^{k}$ and $1\leq j\leq k$ join such pool of nodes. Next, each committee in the remaining butterfly network will “recruit” $b=\frac{3}{4}\log n^{\prime}$ (here $n^{\prime}=\beta n$) random nodes from $\mathcal{V}$. After the recruitment phase, each remaining node $u$ in $\mathcal{V}$ (if any) will probe a random committee in the butterfly. If such a committee has not reached the target size yet it accepts the request, otherwise $u$ tries again with a different committee. It follows that the shrinking phase can be done in $\mathcal{O}(\log n^{\prime})$ rounds w.h.p. where $n^{\prime}=\beta n$.
@@
Let $b=\frac{3n^{\prime}}{4N^{\prime}}=\frac{3}{4}\log n^{\prime}$ and observe that the analysis of the recruitment phase is identical to the one for the growing process. Moreover, after $\Theta(\log n)$ rounds, there can be $n^{\prime}/4$ nodes in $\mathcal{V}$ that are not part of any committee. Each node that joined a committee has one unit of budget that can use to induce a new member into its committee. Every remaining node $u\in\mathcal{V}$ probes a random node $v$ asking to join its committee, if $v$’s budget is greater than 00 then $v$ accepts the request, otherwise $u$ tries again with a different node. Furthermore, observe that even if all the remaining nodes $\mathcal{V}$ found a spot in some committee, there will be $n^{\prime}/2$ nodes that would not have exhausted their budget and each remaining node can find a committee with probability at least $1/2$. Finally, by using similar arguments to the growing process we can show that each remaining node $u$ succeed in finding a committee with high probability in $\Theta(\log n^{\prime})$ rounds. And by using the union bound we can guarantee that each remaining node can find and become part of a committee in $\Theta(\log n^{\prime})$ w.h.p.
1
@@
2The committee $C(0,0)$ is elected as a leader $\ell$.
3Every committee in $\mathcal{S}$ routes its opinion to $\ell$.
4if Agreement on Growing then
5
@@ 6      Each committee $C(i,j)$ for $0\leq i<2^{k}$ and $0\leq j<k$:
15else if Agreement on Shrinking then

### v2

@@ Appendix A Useful Mathematical Tools
$\text{for }R\geq 6\mathbb{E}\left[X\right]$ (1)
@@
$\text{for }k>\mu$ (2)
$\text{for }\varepsilon\geq 0$ (3)
@@
$\text{for }k<\mu$ (4)
$\text{for }\varepsilon\in(0,1)$ (5)
@@ Appendix B Useful Skip list properties
The size of a run of nodes $X_{\ell}$ for some level $\ell$, is at most $\mathcal{O}(\log n)$ w.h.p., and its expected value is 1/(1p)1/(1-p).
@@
The size of a run of nodes $|X_{\ell}|$ is a geometric random variable with parameter 1p1-p. Thus, Choosing $k=c\log_{1-p}n+1$ for $c\geq 1$, gives us a high confidence bound on the number of consecutive nodes in a run of node $X_{\ell}$ at some level $\ell$. Indeed, Moreover, its expected value is $\mathbb{E}\left[|X_{\ell}|\right]=1/(1-p)$, assuming $p=1/2$ (in other words, constant), we have that the expected length of a run of nodes at some level $\ell$ of the skip list is $\mathcal{O}(1)$.
@@
Let $R_{i}$ be number of horizontal edges at level $0\leq i\leq h$ crossed by a search operation that starts on the left topmost node of the skip list. Define the random variable $W_{h}=|R_{0}|+|R_{1}|+\dots+|R_{h}|$ to be the amount of horizontal moves performed by the search algorithm. Observe that each $R_{i}$ is a geometric random variable of parameter 1p1-p and that $h$ is a random variable itself. From Lemma B.1 we know that $\mathbb{E}\left[h\right]=\mathcal{O}(\log n)$ and that . Since $h$ is a random variable, $W_{h}$ is a random sum of random variables thus we can not make a straightforward use of the properties of sum of geometric random variables, rather we can write:
@@
Observe that the first sum is bounded above $16\log n$ because every probability is less than $1$ and the second one is dominated by $\sum_{i\geq 1}1/i^{2}$ which is a constant. Since the running time of a search operation is bounded by the number of horizontal moves performed at each level plus the number of vertical moves to reach the target node from the topmost level to the bottom most. We have that the overall running time of a search procedure starting at the left topmost node in the skip list is $T=h+W_{h}=\mathcal{O}(\log n)$ with probability at least $1-(1/n^{2})$, by applying the union bound we have that $T$ is $\mathcal{O}(\log n)$ with probability at least 1(1/n)1-(1/n) starting at any node in the skip list. To conclude, the expected running time is $\mathbb{E}\left[T\right]=\mathbb{E}\left[h\right]+\mathbb{E}\left[W_{h}\right]=\mathcal{O}(\log n)$.
@@ Appendix C Spartan’s Reshaping protocol
Figure 14: Example of the reshaping procedure. On the left, the Spartan structure as a one-dimensional wrapped butterfly of committees. The rows are represented with their binary representations. Every node in the figure is a random committee of $\Theta(\log n)$ nodes and every edge between two committees encodes a complete bipartite graph. On the right, 22-dimensional butterfly is obtained by increasing the dimensionality of the left-wrapped butterfly by one. Green nodes are the ones that must be added to increase the dimensionality of the $1$-dimensional wrapped butterfly. Moreover, left-to-right execution increases the dimensionality by one, and right-to-left execution decreases the dimensionality by one.
@@
In this section, we provide a detailed description of the reshaping protocol mentioned in Section 2.1. Let $\alpha,\beta\in\mathbb{R}_{\geq 0}$ such that $\alpha>1$ and $\beta\in(0,1)$ be two constants. We want the Spartan network $\mathcal{S}$ to be enlarged if all the committees have size at most $\alpha c\log n$, to shrink if all the committees have size at most $\beta c\log n$, or stay the same otherwise ($c>0$ is the constant hidden by the asymptotic notation). In other words, we want $\mathcal{S}$ to be enlarged or reduced when the wrapped butterfly network size increases or decreases by one dimension respectively. Moreover, a general technique to increase the dimensionality of a $k$-dimensional butterfly network $\mathcal{S}$ by one is to: (1) create a copy of $\mathcal{S}$ in which each row ID $(i,j)$ is “shifted” by $2^{k}$ i.e., $(2^{k}+i,j)$; (2) “shift” the index of the levels of the two disjoint butterflies by one; (3) add a new row of $2^{k+1}$ nodes to form the level 0 of the $k+1$ dimensional butterfly; and, (5) connect the newly created level 0 to the two disjoint butterflies using the “standard” butterfly edge-construction rule [53]. Furthermore, to reduce the dimensionality of a $k$-dimensional butterfly, we need to remove the level 0, remove one of the two disjoint k1k-1-dimensional butterflies, and fix the row and levels IDs if needed. Figure 14 shows an example on how to perform such operations. We start with the description of the growing procedure. A new number of nodes $n^{\prime}=\alpha n$ is reached such that $2^{k}k\in\mathcal{O}(N)$ must be increased to $2^{k+1}(k+1)\in\mathcal{O}(N^{\prime})$ (the number of committees increases from $N=\mathcal{O}(n/\log n)$ to $N^{\prime}=\mathcal{O}(n^{\prime}/\log n^{\prime})$) and update the butterfly structure according to the new size of the network. Moreover, each step of Algorithm 7 requires at most $\mathcal{O}(\log n)$ rounds and no node will send or receive more than $\mathcal{O}(\log n)$ messages at any round. As a first step, the committee $C(0,0)$ is elected as a leader [^16] We can make this assumption because we do not deal with byzantine nodes., and each committee routes a message in which it expresses its opinion (grow, shrink, stay the same) to the leader committee (lines 1-2). This can be done very efficiently on hypercubic networks [53, 61], indeed we have that the leader committee will receive all the opinions in $\mathcal{O}(\log N)$ rounds [54] where $N\in\mathcal{O}(n/\log n)$ is the number of committees in the wrapped butterfly. If the network agrees on growing, the protocol proceeds with increasing the dimensionality of the wrapped butterfly by one. Moreover, assume that the $k$-dimensional butterfly must be transformed into a $k+1$-dimensional one. To this end, the protocol executes the growing approach described before. Every committee shifts its level ID by one and promotes two random nodes in each committee to be the committee leaders of the copy of the butterfly and of the new level 0. Next, all the newly elected committee leaders will start recruiting random nodes in the network until they reach a committee size of $\Theta(\log n^{\prime})$ where $n^{\prime}=\alpha n$. Finally, each new committee will create edges according to the wrapped butterfly construction algorithm [53, 13] and drop the old ones (if any). The next lemma shows that the growing phase requires $\mathcal{O}(\log n^{\prime})$ rounds w.h.p. where $n^{\prime}=\alpha n$.
@@
Let $b=\frac{3n^{\prime}}{4N^{\prime}}=\frac{3}{4}\log n^{\prime}$ we show that after the first phase of the process, each new committee with gain $b$ nodes. During each round, there are at least $\frac{n^{\prime}}{4}-N^{\prime}\in\Omega(n^{\prime})$ nodes that did not receive any invitation from one of these new committee leaders. Moreover, define the event $E_{v}=\text{``The committee leader {v} recruits a node''}$ then setting $N^{\prime}=\frac{n^{\prime}}{c\log n^{\prime}}$ with $c\geq 1$ gives . To bound the expected time needed by a committee to recruit $b$ elements we define a pure-birth Markov Chain $\{X_{t}\}_{t\geq 0}$ with state space $\Omega=\{0,1,2,\dots,b\}$ and initial state 0 that counts the number of recruited members. At each round, the Markov Chain at state $i<b$ can proceed one step forward to state $i+1$ with probability $p_{i,i+1}=$ or loop on $i$ with probability $r_{i,i}=1-p_{i,i+1}$. Observe that each state $0\leq i<b$ is a transient state, while $b$ is an absorbing state, and that the expected absorption time in state $b$ is $\mathcal{O}(b)=\mathcal{O}(\log n^{\prime})$. To give a tail bound on the absorption time, we study a more pessimistic random process in which we toss a p-biased coin with $p=3/16$ and count the number of rounds $Z_{i}$ before we get $b$ heads. In this experiment, the expected number of rounds needed to get $b$ heads is $\frac{16}{3}b\in\Theta(\log n^{\prime})$. By applying a Chernoff bound (Theorem A.1) we can show that for any value of $n^{\prime}$ and any fixed $R$, such that $R\geq 6\mathbb{E}\left[Z\right]$ the probability that we will need more than $R\log n^{\prime}$ rounds to get $b$ number of heads is at most $1/n^{\prime R}$. In other words, this means that the Markov Chain will take at most $\mathcal{O}(\log n^{\prime})$ rounds with probability $1-1/n^{\prime R}$ to reach state $b$. Finally, by applying the union bound we can show that every committee will need $\mathcal{O}(\log n^{\prime})$ rounds with probability $1-1/n^{\prime R-1}$ to recruit $b$ members.
Furthermore, in the case in which the network agreed on shrinking i.e., reducing the butterfly dimensionality from $k$ to k1k-1 a pool of vacated nodes $\mathcal{V}$ is created and all the nodes in the level 0 and in the committees $C(i,j)$ such that $2^{k-1}\leq i<2^{k}$ and $1\leq j\leq k$ join such pool of nodes. Next, each committee in the remaining butterfly network will “recruit” $b=\frac{3}{4}\log n^{\prime}$ (here $n^{\prime}=\beta n$) random nodes from $\mathcal{V}$. After the recruitment phase, each remaining node $u$ in $\mathcal{V}$ (if any) will probe a random committee in the butterfly. If such a committee has not reached the target size yet it accepts the request, otherwise $u$ tries again with a different committee. It follows that the shrinking phase can be done in $\mathcal{O}(\log n^{\prime})$ rounds w.h.p. where $n^{\prime}=\beta n$.
@@
Let $b=\frac{3n^{\prime}}{4N^{\prime}}=\frac{3}{4}\log n^{\prime}$ and observe that the analysis of the recruitment phase is identical to the one for the growing process. Moreover, after $\Theta(\log n)$ rounds, there can be $n^{\prime}/4$ nodes in $\mathcal{V}$ that are not part of any committee. Each node that joined a committee has one unit of budget that can use to induce a new member into its committee. Every remaining node $u\in\mathcal{V}$ probes a random node $v$ asking to join its committee, if $v$’s budget is greater than 0 then $v$ accepts the request, otherwise $u$ tries again with a different node. Furthermore, observe that even if all the remaining nodes $\mathcal{V}$ found a spot in some committee, there will be $n^{\prime}/2$ nodes that would not have exhausted their budget and each remaining node can find a committee with probability at least $1/2$. Finally, by using similar arguments to the growing process we can show that each remaining node $u$ succeed in finding a committee with high probability in $\Theta(\log n^{\prime})$ rounds. And by using the union bound we can guarantee that each remaining node can find and become part of a committee in $\Theta(\log n^{\prime})$ w.h.p.
@@
1The committee $C(0,0)$ is elected as a leader $\ell$.
2Every committee in $\mathcal{S}$ routes its opinion to $\ell$.
3if Agreement on Growing then
4
@@ 5  Each committee $C(i,j)$ for $0\leq i<2^{k}$ and $0\leq j<k$:
14else if Agreement on Shrinking then
@@
Appendix D Detailed execution of the WAVE Protocol
In this section we show a detailed execution of the WAVE protocol. The figures below describe the behavior of the nodes in the clean network $\mathcal{C}$ and the behavior of the nodes in the buffer network $\mathcal{B}$ during the execution of the protocol. We omit the buffer network description when the states of the nodes in $\mathcal{B}$ are clear from the context.
Figure 15: Figure 15(a) shows the Clean Network $\mathcal{C}$ while Figure 15(b) shows the Buffer Network $\mathcal{B}$ after the preproccessing phase. At the beginning, in $\mathcal{B}$, the only cohesive group in the merge state is composed by the nodes in $\mathcal{B}$’s top-most level, with lsls (that is the $\mathcal{B}$’s left sentinel) as a leader. lsls is the node that traverses $\mathcal{C}$ while representing the whole cohesive group $X=\{ls,23,98,rs\}$.
Figure 16: The unique cohesive group $X$ splits into two cohesive groups: $X_{\texttt{down}}=\{ls\}$ and $X_{\texttt{right}}=\{23,98,rs\}$. $X_{\texttt{down}}$ merges in level $3$ and proceeds downwards while $X_{\texttt{right}}$ (lead by the new $23$ as the new leader) moves right to find its designated spot to merge at level $3$. Every step executed by $X_{\texttt{down}}$ and $X_{\texttt{right}}$ on $\mathcal{C}$ is communicated to their respective children that, upon receiving new information, update their anchor points (i.e., their virtual walk position).
Figure 17: $X_{\texttt{down}}=\{ls\}$ keeps merging at level 22 while $X_{\texttt{right}}=\{23,98,rs\}$ merges at level 33 between $13$ and $\infty$ and moves one step down. Their children update their virtual walk status as well.
Figure 18: $X_{\texttt{down}}=\{ls\}$ completed the merging in $\mathcal{C}$, consequently the node 11 starts its own merge phase at lsls in level 0 and in $\mathcal{O}(1)$ rounds finds its spot in the base level of $\mathcal{C}$ completing its own merge phase. $X_{\texttt{right}}=\{23,98,rs\}$ splits in two, $X_{\texttt{down}}^{\prime}=\{23\}$ and $X_{\texttt{right}}^{\prime}=\{98,rs\}$ where $X_{\texttt{down}}^{\prime}$ continues merging downwards while $X_{\texttt{right}}^{\prime}$ moves right in level $1$. Finally, node $25$ starts its merge phase at level 11 starting from $23$ in $\mathcal{C}$, and, being dependent on $23$, waits for $25$ to merge in level 0 before proceeding.
Figure 19: $X_{\texttt{down}}^{\prime}=\{23\}$ complete its own merge phase by merging at level level 0. Consequently, $25$ merges in level 0 in $\mathcal{O}(1)$ rounds. On the other side of the skip list, $X_{\texttt{right}}^{\prime}=\{98,rs\}$ merges at level $1$ and jumps one level below, while 55 “virtually” waits for its turn at node $50$ in $\mathcal{C}$.
Figure 20: $X_{\texttt{right}}^{\prime}=\{98,rs\}$ completes its merge phase by merging in the bottom most level, 55 enters the merge phase starting at $50$ and successfully merges in $\mathcal{O}(1)$ rounds in level 0.